OpenStudy (anonymous):

You and a friend bought the same number of CDs and the same number of DVDs at two different stores. Your store charges $6 per CD, $9 per DVD, and you paid $36. Your friends store charges $7 per CD, $8 per DVD, and he paid $37. How many CDs did you buy?

OpenStudy (anonymous):

set up two equations 6x+9y=36 7x+8y=37

OpenStudy (anonymous):

use one equation to solve for one variable in terms of the other, i solve for x using the first euqation 6x=36-9y x=6-3/2y

OpenStudy (anonymous):

plug this into the other equation 7(6-3/2y)+8y=37 42-21/2y+8y=37 (-21/2+16/2)*y=-5 -5/2*y=-5 y=2

OpenStudy (anonymous):

now plug this into the original equations 6*x+9*2=36 6*x+18=36 6*x=18 x=3

OpenStudy (anonymous):

now try these in the other equation to test your answer 7*3+8*2 21+16 37

OpenStudy (anonymous):

so the answer is 3 cd's, 2 dvd's